The students of two groups A and B obtained following marks in a test of Mathematics.


Find which group gave more consistent performance, based on coefficient of range.


Marks of Group A 35 47 42 45 40 39 38 46 41 43


Marks of Group B 35 70 55 65 60 42 73 80 53 57


4. The temperatures recorded at the end of every 4 hours during a day at two cities Mumbai


and Pune are given below:


Mumbai 28 30 34 31 28 27


Pune 25 28 34 40 32 26


Use appropriate measure of dispersion and comment on variation.

1
Expert's answer
2022-01-10T18:00:35-0500

3.

Coefficient of Range:

"k=\frac{f_{max}-f_{min}}{f_{max}+f_{min}}"


for Group A:

"k_1=\frac{47-35}{47+35}=0.146"


for Group B:

"k_2=\frac{80-35}{80+35}=0.391"


since "k_1<k_2" , Group A has more consistent performance


4.

standard deviation:

"s=\sqrt{\frac{\sum (x_i-\overline{x})^2}{n-1}}"


for Mumbai:

"s_1=2.582"


for Pune:

"s_2=5.672"


since "s_1<s_2" , we can conclude that dispersion of temperature around mean value

is more for Pune